# Star formation rate density as a function of galaxy mass at z 0.2 with MUSE and GAMA surveys

### v1

@@ 3 Derivation of the Star Formation Rate Density with the method
3.1 Star Formation Rate from emission lines
Star forming galaxies produce very specific emission lines in their spectra that can be easily characterised. emission lines are one of the best indicators for star forming galaxies, since only young, massive stars with ages 20 Myr produce vast amounts of photons that ionise the surrounding gas (Davies et al., 2016). This is a great advantage when calculating the SFR of a galaxy, as only young stars are taken into account and thus it is largely independent of SF history (Kennicutt Jr, 1998). Assuming a Kroupa initial mass function (Kroupa, 2001), the relation between the luminosity and star formation rate may be obtained with the following calibration (Calzetti, 2013):
with in M yr-1 and in erg s-1. The calibration constant is valid provided that star formation has remained constant over timescales 6 Myr, with no dependency on long timescales (Calzetti, 2013). This calibration does not take into account AGN contamination, which could theoretically result in an overestimation of SFR(H). However, for low-mass galaxies AGN contamination is not a significant issue Kauffmann et al. (2003)
@@ 3.2 Star Formation Rate Density with the method
In order to calculate the SFRD with respect to mass, it is first necessary to account for the Malmquist bias (Schmidt, 1968). The Malmquist bias is a selection bias that applies to astronomical surveys where intrinsically bright objects are favoured over dimmer ones. Dimmer objects are not bright enough to be detected at further distances, but are only observed at closer ones. On the other hand, bright objects are detected at nearer and further distances, causing a strong statistical bias towards brighter objects. To counteract this bias, one of the most common correction method is to use is the method (Felten, 1976), where indicates the maximum volume over which the objects can be observed given their luminosity and within the chosen redshift limit. The method ensures that brighter objects, as they have a larger maximum volume, have less statistical weight. The comoving volume may be derived using the following equation:
where is the solid angle of the FOV of the survey, and and are the maximum and minimum comoving distance for which the object could be observed given its luminosity and survey limits.
@@ 3.3 Estimation of
In order to correctly apply the 1/Vmax method described in Sec. 3.2, it is necessary to estimate the maximum distance at which a source can be observed given its luminosity. This can be done in different ways. For the sample from MUSE WIDE and MUSE HUDF, Dmax was derived by defining a minimum flux detection limit (see Sec. 4.2 for derivation). If the detection limit is defined as the minimum flux that a source must have in order to be included in the sample (), one can work out the maximum luminosity distance () at which it can be observed by rearranging the relationship between bolometric luminosity and flux:
In the case of bolometric flux, i.e. with no k-correction, then
@@
where is the k-correction of the observed galaxy for the effective band at zmax (see Blanton et al. 2003 for this nomenclature). This is obtained by iterating: starting from and determining a new by interpolation from a vector of band-shifted k-correction values on each iteration. Note that , ensuring that converges to zero when .
@@ 3.4 Dust extinction
In order to accurately measure SFR, dust extinction corrections must be taken into account. H line is situated in the optical spectrum, making it less sensitive to dust extinction compared to other star formation indicators such as UV light. However, even for local galaxies, dust effects on H is significant (Koyama et al., 2015). A common practice to determine the levels of dust attenuation in galaxies is to measure the ratio between H and H, also known as the Balmer decrement. The theoretical value of the Balmer decrement, assuming that a optically thick star forming region for all Lyman lines greater than Ly, is equal to 2.86 (Calzetti et al., 2000). Therefore, for a fixed electron temperature of K, one can assume that any deviation from the theoretical value is due to dust extinction (Groves et al., 2012).
The magnitude extinction from the observed Balmer decrement may be calculated using the relation described in Osterbrock & Ferland (2006), with the E(B-V) colour excess that can be calculated directly from the Balmer decrement using the equation (see (Domínguez et al., 2013)):
with the extinction coefficients and coming from the reddening curve between H and H described in Cardelli et al. (1989). The dust correction factor is then derived with Calzetti et al. (2000):
@@ 4.1 Sample selection
Initially, all the galaxies in the MUSE Wide and HUDF catalogues (Urrutia et al., 2019; Bacon et al., 2023) that presented emission lines at redshift 0.2 were considered. Since the two areas covered by HUDF and WIDE overlay slightly, it was also ensured that there was no duplicated galaxies, resulting in a preliminary sample of 21 star forming galaxies from MUSE Wide emission line catalogue and 10 from MUSE HUDF main catalogue. In order to verify the accuracy of the redshift measurements, AUTOZ, which is a fully automated redshift code that allows to homogenise redshift measurements (Baldry et al., 2014), was run over the entire area. The test resulted in no significant difference between the values reported in the MUSE Wide and HUDF catalogues, and the redshifts derived using AUTOZ.
MUSE Wide DR1 and MUSE HUDF DR2 used different photometry to derive the stellar masses values for the galaxies, the Skelton et al. (2014) and the HST R15 Rafelski et al. (2015) catalogues respectively. The HST R15 Rafelski et al. (2015) catalogue used in HUDF, does not contain values for F160W magnitude for all the sources in the catalogue. Therefore, photometry from the van der Wel et al. (2014) catalogue was used instead, as it contained values for all the sources in the sample. In order to ensure consistency, the photometric catalogues the F160W magnitudes were compared, showing no notable distinction, as it can be seen in the scatter plot in Fig. 1. This allowed us to assume that using photometric values from either catalogue does not affect the derivation of the H-band luminosity of the sources.
Using the F160W magnitude values from the van der Wel et al. (2014) catalogue, the H-band H luminosities of all the galaxies in the sample were calculated. This information was then used to derive the stellar mass of two sources in the HUDF region that do not have a valid derivation for the stellar mass using MagPhys. In fact, Source ID6474 does not have a photometric value in R15, whereas the stellar mass of source ID8051 is likely extremely low. In order to include them in the sample, as they both present emission lines, their stellar masses were derived using the typical stellar mass-to-luminosity ratio of star forming galaxies.
@@ 4.2 Detection Limit
Estimating the detection limits of the surveys is necessary in order to perform the the method (Schmidt, 1968). In fact, because they affect the maximum volume within which objects can be detected, a failure to account for this bias can lead to incorrect estimates of the space density of the population being studied. However, the derivation of detection limits is often very complex, as there are many factors such as sky lines, atmospheric conditions and aperture sizes, that need to be considered.
In order to simplify the derivation of the detection limit and the calculation of Vmax, an empirical detection limit (i.e. sample-selection limit) was set by analysing all the emission line fluxes detected in the survey between 6563 and 7876 Å, which is the region where is expected to be found at . A reasonable assumption would be that the highest number of observations will have values just above the detection limit, and that below the limit there will be a sharp decrease in detections, since the instrument would have more difficulty in detecting the signal below a certain threshold.
Using the histogram shown in Fig. 4 and 5, the detection limit was estimated. It can be seen how there is a significant decline in number of emission lines observed right before the peak in the histogram. The detection limits were thus set to be the cut-off values where the number of the detections dropped, which can be seen in the plots as the red dashed line. The plots also show all the emission line fluxes detected between 6563 and 7876 Å, with the red points representing the lines. The results show that all the emissions in MUSE Wide, and all but one in MUSE HUDF, are above the empirically estimated detection limits. The source that is below this limit (MUSE HUDF ID8061) is excluded from the estimation of the SFRD.
@@ 4.3 Dust effects
Since H is affected by dust extinction in the optical range, it is essential to consider dust effects on emission line flux measurements in order to derive accurate SFR results. As described in Sec. 3.4, the most robust way to estimate dust extinction in a galaxy is by measuring the flux ratio between H and H, also known as the Balmer decrement. One caveat of this method is that the H line is often very faint and undetectable, particularly for smaller, fainter galaxies such as the sources in our sample. To overcome the absence of H lines in our sample, the Balmer decrement for galaxies of similar properties was derived from the Sloan Digital Sky Survey (SDSS) (York et al., 2000) and used to estimate the typical Balmer decrement of star-forming galaxies in a mass bin.
From SDSS, all the galaxies with valid H and H emission line fluxes and H S/N 20 were considered. Then, the Balmer decrement for each galaxy was calculated, before identifying the median value for each stellar mass bin of M = 0.5. Using the median value for each mass bin, the relation between Balmer decrement and stellar mass was derived. This relation has been observed many times over the past decades, showing that the Balmer decrement, and thus dust, is higher in higher mass galaxies (Domínguez et al., 2013; Garn & Best, 2010). The results are reported in Fig. 6, where the red exponential line shows the relation between Balmer decrement of star forming galaxies and their stellar mass, where it is possible to see how galaxies with lower stellar masses are significantly less affected by dust. The resulting values for Balmer decrement and dust extinction for each mass bin are reported in Table 1.
| Mass bin | Balmer Decrement |
| 5.5 - 6.0 | 2.980136 |
| 6.0 - 6.5 | 2.993040 |
| 6.5 - 7.0 | 3.012704 |
| 7.0 - 7.5 | 3.042670 |
| 7.5 - 8.0 | 3.088333 |
| 8.0 - 8.5 | 3.157919 |
| 8.5 - 9.0 | 3.263959 |
| 9.0 - 9.5 | 3.425550 |
@@ 4.4 H-alpha luminosity function
In order to place the depth of the H-alpha selected sample into context, we compute the H luminosity function using the values for the combined MUSE sample. The observed luminosity function (i.e. not corrected for dust) is shown in the plot in Fig. 7, where it is compared with the results obtained by Gilbank et al. (2010b). This shows that the MUSE sample is probing to significantly fainter luminosities than the SDSS sample of Gilbank et al. The latter sample was restricted to . The SDSS spectra are reasonably well calibrated and the aperture corrections were made using the -band in the higher-S/N SDSS Stripe 82 region.
For the GAMA sample, estimating H luminosity functions is significantly harder because of the smaller aperture used for the majority of spectra and the less accurate spectrophotometry (see Hopkins et al. 2013 for GAMA spectroscopic analysis). Gunawardhana et al. (2013) compute dust-corrected H luminosity functions for a similar GAMA sample and compare with other surveys as well. They discuss the problem of bivariate selection and the discrepancies between various results. We note the MUSE sample, while H selected, is over a small volume and is likely underdense as will be discussed in the § 6.
@@ 5 Results: the SFRD as a function of mass
From the dust corrected line fluxes measured by MUSE Wide and MUSE HUDF, the SFRs of the galaxies were calculated using the calibration shown in Eq. 1. Two star-forming galaxies, one from MUSE Wide and one from MUSE HUDF, although they have emission lines, do not present any confirmed emission, so they were excluded from the analysis, resulting in a total of 27 sources (19 from Wide and 8 from from HUDF). Before deriving the SFRD, the maximum volume of the survey was calculated using the detection limits determined previously. Fig. 8 shows the maximum volume of each galaxy, with derived considering the total volume of the two surveys combined, as described in sec 3.2. The plot shows that the galaxies with the lowest stellar masses tend to have lower , which is consistent with the expectation, as they tend to have lower H flux and thus, they would not be observed at further distances.
@@
The SFRD for the combined MUSE Wide and MUSE HUDF sample was derived using the method with Eq. 4 described in Sec. 3.2. Given the limited number of galaxies in the sample, the mass bins were chosen to be relatively wide, with a value of . This was necessary in order to have a reasonable estimate of the uncertainty using pseudo-Poisson weighted errors for each bin (sum of weights squared). The SFRD was calculated starting from two different mass bins, and , resulting in two overlapping functions for the same sample.
@@
The SFRD was also calculated using a sample from the GAMA survey, a large magnitude-limited spectroscopic survey. Compared to MUSE Wide and MUSE HUDF, GAMA is shallower (and missing low-surface-brightness galaxies) but with a much larger area, comprising a vaster number of galaxies and thus also providing a good indication of the number density at lower stellar masses. The GAMA catalogue used contained measurements of spectroscopic redshifts from the SpecCat DMU, stellar masses from StellarMasses DMU, and SFR measurements at different timescales derived through SED fitting with MagPhys in the MagPhys DMU. SED measurements of SFRs were used instead of deriving the SFR using measurements of H emission line flux measurements, as GAMA fiber spectroscopy may overestimate their values for dwarf galaxies (Richards et al., 2016). Since only young, massive stars with ages 20 Myr produce H emissions (Davies et al., 2016), the median SFR at timescales of 107 and 108 were used in this analysis (from the MagPhys table). After ensuring that only galaxies with valid measurements were included, the final sample from GAMA consisted of 7579 galaxies at redshift z 0.06 from the equatorial regions that are highly complete (179 deg2).
@@
The results for the SFRD using the combined MUSE sample at different mass bins and using the GAMA sample at different timescales can be seen in Fig. 9. The plot shows that the samples appear to concur, with a steady decline at the lower mass end of the SFRD(M). The decline may be quantified by assuming a linear regression between SFRD and stellar mass in log space at lower stellar masses, and measuring its slope , i.e.
@@
Considering the relation between SFR and stellar mass in star-forming galaxies, the SFR of a galaxy is approximately proportional to its stellar mass so that
with 1, at masses below about 1010 , along the star-forming main sequence (Brinchmann et al., 2004; Lee et al., 2015; McGaugh et al., 2017). Since the SFRD is the number density times the mean SFR (e.g. Behroozi et al. 2013), the relationship between the slope of the SFRD and GSMF can be expressed as:
with along the main sequence, the equation becomes: . Therefore, the values of the slope of the SFRD implies in the range to . These values are consistent with previous studies of the GSMF (Moustakas et al., 2013; Wright et al., 2017; Thorne et al., 2021; Driver et al., 2022), providing further evidence of no turnover at lower masses. This would mean that the number density of low-mass galaxies is high, contributing significantly to the star formation activity without any indication of flattening or declining.

### v2

@@ 3 Derivation of the Star Formation Rate Density with the method
3.1 Star Formation Rate from H emission lines
Star forming galaxies produce very specific emission lines in their spectra that can be easily characterised. H emission lines are one of the best indicators for star forming galaxies, since only young, massive stars with ages 20 Myr produce vast amounts of photons that ionise the surrounding gas (Davies etĀ al., 2016). This is a great advantage when calculating the SFR of a galaxy, as only young stars are taken into account and thus it is largely independent of SF history (Kennicutt, 1998). Assuming a Kroupa initial mass function (Kroupa, 2001), the relation between the H luminosity and star formation rate may be obtained with the following calibration (Calzetti, 2013):
with SFR in M yr-1 and in erg s-1. The calibration constant is valid provided that star formation has remained constant over timescales 6 Myr, with no dependency on long timescales (Calzetti, 2013). This calibration does not take into account AGN contamination, which could theoretically result in an overestimation of SFR(H). However, for low-mass galaxies AGN contamination is not a significant issue Kauffmann etĀ al. (2003)
@@ 3.2 Star Formation Rate Density with the method
The SFRD with respect to mass can be calculated using the method. The method (Felten, 1976) corrects for survey biases by weighting galaxies according to the maximum comoving volume in which they could be detected. This ensures that fainter galaxies, observable over smaller volumes, contribute proportionately more to the overall density estimate. The comoving volume may be derived using the following equation:
where is the solid angle of the FOV of the survey, and and are the maximum and minimum comoving distance for which the object could be observed given its luminosity and considering the survey limits. Note for a redshift-limited sample, the brighter sources that are limited only by the redshift range will have the same .
@@ 3.3 Estimation of
In order to correctly apply the 1/Vmax method described in Sec.Ā 3.2, it is necessary to estimate the maximum distance at which a source can be observed given its luminosity. This can be done in different ways. For the sample from MUSE WIDE and MUSE HUDF, Dmax was derived by defining a minimum flux detection limit (see Sec. 4.2 for derivation). If the detection limit is defined as the minimum flux that a source must have in order to be included in the sample (), one can work out the maximum luminosity distance () at which it can be observed by rearranging the relationship between bolometric luminosity and flux:
In the case of bolometric flux, i.e.Ā with no k-correction, then
@@
where is the k-correction of the observed galaxy for the effective band at (see Blanton etĀ al. 2003 for this nomenclature). This is obtained by iterating: starting from and determining a new by interpolation from a vector of band-shifted k-correction values on each iteration. Note that , ensuring that converges to zero when .
@@ 3.4 Dust extinction
In order to accurately measure the SFR, dust extinction corrections must be taken into account. H line is situated in the optical spectrum, making it less sensitive to dust extinction compared to other star formation indicators such as UV light. However, even for local galaxies, dust effects on H are significant (Koyama etĀ al., 2015). A common practice to determine the levels of dust attenuation in galaxies is to measure the ratio between H and H, also known as the Balmer decrement. The theoretical value of the Balmer decrement, assuming that a optically thick star forming region for all Lyman lines greater than Ly, is equal to 2.86 (Calzetti etĀ al., 2000). Therefore, for a fixed electron temperature of K, one can assume that any deviation from the theoretical value is due to dust extinction (Groves etĀ al., 2012).
The magnitude extinction from the observed Balmer decrement may be calculated using the relation described in Osterbrock & Ferland (2006), with the E(B-V) colour excess that can be calculated directly from the Balmer decrement using the equation (see (DomĆnguez etĀ al., 2013)):
with the extinction coefficients and coming from the reddening curve between H and H described in Cardelli etĀ al. (1989). The dust correction factor is then derived with Calzetti etĀ al. (2000):
@@ 4.1 Sample selection
Initially, all the galaxies in the MUSE Wide and HUDF catalogues (Urrutia etĀ al., 2019; Bacon etĀ al., 2023) that presented H emission lines at redshift 0.2 were considered. Since the two areas covered by HUDF and WIDE overlay slightly, it was also ensured that there were no duplicated galaxies, resulting in a preliminary sample of 21 star-forming galaxies from MUSE Wide emission line catalogue and 10 from MUSE HUDF main catalogue. In order to verify the accuracy of the redshift measurements, AUTOZ, which is a fully automated redshift code that allows us to homogenise redshift measurements (Baldry etĀ al., 2014), was run over the entire area. The test resulted in no significant difference between the values reported in the MUSE Wide and HUDF catalogues, and the redshifts derived using AUTOZ.
MUSE Wide DR1 and MUSE HUDF DR2 used different photometry to derive the stellar masses values for the galaxies, the Skelton etĀ al. (2014) and the HST R15 Rafelski etĀ al. (2015) catalogues respectively. The HST R15 Rafelski etĀ al. (2015) catalogue used in HUDF, does not contain values for F160W magnitude for all the sources in the catalogue. Therefore, photometry from the vanĀ der Wel etĀ al. (2014) catalogue was used instead, as it contained values for all the sources in the sample. In order to ensure consistency, the photometric catalogues the F160W magnitudes were compared. This allowed us to assume that using photometric values from either catalogue does not affect the derivation of the H-band luminosity of the sources.
Using the F160W magnitude values from the vanĀ der Wel etĀ al. (2014) catalogue, the H-band luminosities of all the galaxies in the sample were calculated. This information was then used to derive the stellar mass of two sources in the HUDF region that do not have a valid derivation for the stellar mass using MagPhys. In fact, Source ID6474 does not have a photometric value in R15, whereas the stellar mass of source ID8051 is likely extremely low. In order to include them in the sample, as they both present H emission lines, their stellar masses were derived using the typical stellar mass-to-luminosity ratio of star forming galaxies.
@@ 4.2 Detection Limit
Estimating the detection limits of the surveys is necessary in order to perform the method (Schmidt, 1968). In fact, because they affect the maximum volume within which objects can be detected, a failure to account for this bias can lead to incorrect estimates of the space density of the population being studied. However, the derivation of detection limits is often very complex, as there are many factors such as sky lines, atmospheric conditions, and aperture sizes that need to be considered.
In order to simplify the derivation of the detection limit and the calculation of Vmax, an empirical detection limit (i.e.Ā sample-selection limit) was set by analysing all the emission line fluxes detected in the survey between 6563 and 7876 Ć , which is the region where H is expected to be found at . A reasonable assumption would be that the highest number of observations will have values just above the detection limit, and that below the limit there will be a sharp decrease in detections, since the instrument would have more difficulty in detecting the signal below a certain threshold.
Using the histogram shown in Fig. 3, the detection limit was estimated. It can be seen that there is a significant decline in number of emission lines observed right before the peak in the histogram. The detection limits were thus set to be the cut-off values where the number of detections dropped, which can be seen in the plots as the red dashed line. The plots also show all the emission line fluxes detected between 6563 and 7876 Ć , with the red points representing the H lines. The results show that all the H emissions in MUSE Wide, and all but one in MUSE HUDF, are above the empirically estimated detection limits. The source below this limit (MUSE HUDF ID8061) is excluded from the estimation of the SFRD.
@@ 4.3 Dust effects
Since H is affected by dust extinction in the optical range, it is essential to consider dust effects on emission line flux measurements in order to derive accurate SFR results. As described in Sec. 3.4, the most robust way to estimate dust extinction in a galaxy is by measuring the flux ratio between H and H, also known as the Balmer decrement. One caveat of this method is that the H line is often very faint and undetectable, particularly for smaller, fainter galaxies such as the sources in our sample. To overcome the absence of H lines in our sample, the Balmer decrement for galaxies of similar properties was derived from the Sloan Digital Sky Survey (SDSS) (York etĀ al., 2000) and used to estimate the typical Balmer decrement of star-forming galaxies in a mass bin. This approach assumes that the median Balmer decrement trend with stellar mass also applies to the MUSE sample.
The total stellar mass of star-forming galaxies correlates with average metallicity, size and dust mass (Beeston etĀ al., 2018). So naturally, the dust extinction is lower in low-mass galaxies because less sight lines go through high extinction regions. Thus for our purposes, where we do not have individual Balmer decrement measurements, an average value as a function of stellar mass is the best estimate we can make.
From SDSS, all the galaxies with measured H and H emission line fluxes (Brinchmann etĀ al., 2004; Tremonti etĀ al., 2004) and H S/N 20 were considered. Then, the Balmer decrement for each galaxy was calculated, before determining the median value for each stellar mass bin of M = 0.5. From these points, a simple relation between Balmer decrement and stellar mass was derived using a quadratic function for stellar masses larger than , and assuming a constant decrement at lower masses. This provides a robust estimate that is sufficient for the scope of this study. The results are reported in Fig.Ā 4, where the red line shows the relation between Balmer decrement of star forming galaxies and their stellar mass, where it is possible to see how galaxies with lower stellar masses are significantly less affected by dust. The resulting values for Balmer decrement and dust extinction for each mass bin are reported in TableĀ 1.
This dust-extinction-mass relation has been observed many times over the past decades, showing that the Balmer decrement, and thus dust extinction, is significantly larger in high mass galaxies (DomĆnguez etĀ al., 2013; Garn & Best, 2010). For low-mass galaxies, the accuracy of the SFRD is not limited by the dust extinction estimate (only a correction factor of 1.2) whereas for high-mass galaxies (), it is arguably the limiting factor.
| Mass bin | Balmer Decrement | Dust correction factor |
| 3.07 | 1.20 | |
| 7.5 ā 8.0 | 3.07 | 1.20 |
| 8.0 ā 8.5 | 3.09 | 1.22 |
| 8.5 ā 9.0 | 3.20 | 1.35 |
| 9.0 ā 9.5 | 3.41 | 1.58 |
@@ 4.4 H-alpha luminosity function
In order to place the depth of the H-alpha selected sample into context, we compute the H luminosity function using the values for the combined MUSE sample. The observed H luminosity function (i.e. not corrected for dust) is shown in the plot in Fig. 5, where it is compared with the results obtained by Gilbank et al. (2010b). This shows that the MUSE sample is probing to significantly fainter luminosities than the SDSS sample of Gilbank et al. The latter sample was restricted to . The SDSS spectra are reasonably well calibrated and the aperture corrections were made using the -band in the higher-S/N SDSS Stripe 82 region. The offset evident in the figure is likely due to the underdensity of the MUSE sample discussed in § 6.
For the GAMA sample, estimating H luminosity functions is significantly harder because of the smaller aperture used for the majority of spectra and the less accurate spectrophotometry (see Hopkins etĀ al. 2013 for GAMA spectroscopic analysis). Gunawardhana etĀ al. (2013) compute dust-corrected H luminosity functions for a similar GAMA sample and compare with other surveys as well. They discuss the problem of bivariate selection and the discrepancies between various results.
@@ 5 Results: the SFRD as a function of mass
From the dust corrected H line fluxes measured by MUSE Wide and MUSE HUDF, the SFRs of the galaxies were calculated using the calibration shown in Eq.Ā 1. Two star-forming galaxies, one from MUSE Wide and one from MUSE HUDF, although they have emission lines, do not present any confirmed H emission, so they were excluded from the analysis, resulting in a total of 27 sources (19 from Wide and 8 from from HUDF). Before deriving the SFRD, the maximum volume of the survey was calculated using the detection limits determined previously. Fig. 6 shows the maximum volume of each galaxy, with derived considering the total volume of the two surveys combined, as described in sec 3.2. The plot shows that the galaxies with the lowest stellar masses tend to have lower , which is consistent with the expectation, as they tend to have lower H flux and thus, they would not be observed at further distances.
@@
The SFRD for the combined MUSE Wide and MUSE HUDF sample was derived using the method with Eq. 4 described in Sec. 3.2. Given the limited number of galaxies in the sample, the mass bins were chosen to be relatively wide, with a value of . This was necessary in order to have a reasonable estimate of the uncertainty using weighted-Poisson errors for each bin (Bohm & Zech 2014; the variance is the sum of weights squared with the weights as per Eq.Ā 4). The SFRD was calculated starting from two different mass bins, and , resulting in two overlapping functions for the same sample.
@@
The SFRD was also calculated using a sample from the GAMA survey, a large magnitude-limited spectroscopic survey. Compared to MUSE Wide and MUSE HUDF, GAMA is shallower (and missing low-surface-brightness galaxies) but with a much larger area, comprising a vaster number of galaxies and thus also providing a good indication of the number density at lower stellar masses. The GAMA catalogue used contained measurements of spectroscopic redshifts from the SpecCat DMU, stellar masses from StellarMasses DMU, and SFR measurements at different timescales derived through SED fitting with MagPhys in the MagPhys DMU. SED measurements of SFRs were used instead of deriving the SFR using measurements of H emission line flux measurements, as GAMA fiber spectroscopy may overestimate their values for dwarf galaxies (Richards etĀ al., 2016). Since only young, massive stars with ages 20 Myr produce H emissions (Davies etĀ al., 2016), the median SFR at timescales of 107 and 108 were used in this analysis (from the MagPhys table). After ensuring that only galaxies with valid measurements were included, the final sample from GAMA consisted of 7579 galaxies at redshift from the equatorial regions that are highly complete (179 deg2).
@@
The results for the SFRD using the combined MUSE sample at different mass bins and using the GAMA sample at different timescales can be seen in Fig. 7. The plot shows that the samples appear to concur, with a steady decline at the lower mass end of the SFRD(). The decline may be quantified by assuming a linear regression between SFRD and stellar mass in log space at lower stellar masses, and measuring its slope , i.e.
@@
Considering the relation between SFR and stellar mass of star-forming galaxies, the SFR of a galaxy is approximately proportional to its stellar mass so that
with , at masses below about , along the star-forming main sequence (Brinchmann etĀ al., 2004; Lee etĀ al., 2015; McGaugh etĀ al., 2017). Since the SFRD is the number density times the mean SFR (e.g.Ā Behroozi etĀ al. 2013), the relationship between the slopes of the SFRD, SFR-mass relation and GSMF (of star-forming galaxies) can be expressed as:
This is accurate over any mass range where there is no significant break in the power laws.
It is possible to estimate the GSMF directly from these samples, however, they are expected to be incomplete even for star-forming galaxies at low masses. This especially true for H selection which is highly biased toward high-SFR galaxies, with significantly reduced detectability for low SFRs (lower luminosity, equivalent width and surface brightness). Instead if we assume along the main sequence of star-forming galaxies, the above relationship (Eq.Ā 15) becomes . Therefore, the values of the slope of the SFRD imply in the range to . These values are consistent with previous studies of the GSMF (Moustakas etĀ al., 2013; Wright etĀ al., 2017; Thorne etĀ al., 2021; Driver etĀ al., 2022), providing further evidence of no turnover at lower masses.